# Domain Adaptation from Generated Multi-Weather Images for Unsupervised Maritime Object Classification

### v2

@@ 3.2. Generalization Enhancement Based on CLIP
VLMs exhibit exceptional performance across various distributions, attributed to the vast diversity of distributions seen during their training (Radford et al., 2021), which leverage contrastive pre-training methods (Zhai et al., 2022) for learning the correlation between the embeddings of matching image-text pairs. For instance, CLIP (Radford et al., 2021) are trained jointly on 400 million image-text pairs, yielding remarkable generalization across different data distributions. Despite the considerable zero-shot generalization of VLMs in classification tasks, effectively adapting them to downstream tasks remains a great challenge. It is expected to leverage the knowledge of VLM in a portable way. Knowledge distillation not only avoids large-scale parameter tuning of VLMs, but also helps to realize lightweight reasoning. For UDA task, it is an effective strategy that enhances cross-domain representation across the open semantic and visual space of VLMs, in order to bridge the domain gap. Compared to a single label, VLMs have more descriptive information, which is beneficial of making full use of the AIMO’s rich label information as semantic consistency constraint. Therefore, we propose a Generalization Enhancement module to efficiently distill the abundant knowledge from CLIP into the adversarial UDA network. We distill the knowledge of text embeddings and image embeddings extracted by frozen VLMs into feature extractor of ViT to enhance feature representations, by carrying out a two-branch constraint to simultaneously align backbone features with semantic information and visual information from VLMs.
@@
Specifically, we perform a Self-Knowledge Distillation on the source domain data based on CLIP. The feature invariance is constructed in the source domain by Self-Knowledge Distillation, which makes the model more robust to the distribution migration in the target domain. Firstly, We utilize the frozen text encoder and image encoder of VLMs to extract embedded features, which update the source domain image representations as a ”teacher signal”. And then the similarity between text embeddings, image embeddings and backbone features is calculated respectively, as constraints to supervise the training process, in order to realize a double alignment of semantic space and visual space. Given labels for category and weather conditions, as well as images from the source domain as inputs to CLIP, the source domain features of Transformer Blocks are aligned with the text-encoder and image-encoder features of CLIP in the Generalization Enhancement module. As proper text prompt allow for better control of invariance in text embeddings (Addepalli et al., 2024), we use ”A photo of {class} in {domain} to represent the semantic embedding. {domain} is referred to the weather and illumination condition of the source image. We denote the CLIP’s text embedding for ”A photo of {class} in {domain} and image embedding of the source domain input $(\bm{x}^{s}_{i},\bm{y}^{s}_{i})$ as $\bm{T}_{{y}^{s}_{i}}$ and $\bm{I}_{\bm{x}^{s}_{i}}$. The generalization enhanced feature corresponding to the input $\bm{x}^{s}_{i}$ are denoted as $\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}_{i}}$. The following Self-Knowledge Distillation loss is used to refine the aligned features:
(4) $\mathcal{L}_{SKD}=-\mathbb{E}_{(\bm{x},\bm{y})\in\mathcal{D}_{s}}\sum_{i=1}^{n_{s}}\{\cos(\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}_{i}},\bm{T}_{{y}^{s}_{i}})+\cos(\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}_{i}},\bm{I}_{\bm{x}^{s}_{i}})\}.$
@@
3.3. Classifier Refinement
We regularize the latent feature spaces of the transformer backbone through classifier refinement using perturbed source and target domain features to reduce overfitting . Incorporating a random offset into the data is a prevalent data augmentation technique that is widely employed to alleviate overfitting. It may be better to perform data augmentation operations at hidden layers rather than directly manipulating the input images (Verma et al., 2019). The optimal layer for introducing perturbations differs from various tasks. Perturbing relatively deep layers tends to yield superior performance but increases the risk of model collapse, and randomly choosing one transformer block proves to be more robust compared with perturbing any individual layer from them (Sun et al., 2022a). Therefore, we select a transformer block at random, and then add a random offset to the input token sequence of the chosen block.
@@ 3.3. Classifier Refinement
(6) $\bm{D}_{\text{KL}}(\bm{p}||\bm{\tilde{p}})=\sum_{i=1}\bm{p}(i)\log\frac{\bm{p}(i)}{\bm{\tilde{p}}(i)}.$
$\bm{D}_{\text{KL}}(\bm{\tilde{p}}||\bm{p})$ is the calculated similarly. And it is more robust to combine the two KL divergence together in random offset loss (Sun et al., 2022a). To achieve reliable predictions, we define a Confidence Filter $\mathcal{F}$ as
(7) $\mathcal{F}[\mathcal{D};\bm{p}]=\{\bm{x}\in\mathcal{D}|\max(\bm{p})>\kappa\},$
@@
(8) $\begin{split}\mathcal{L}_{offset}=\mathbb{E}_{\mathcal{B}\subseteq\mathcal{D}_{\bm{s}}\cup\mathcal{D}_{\bm{t}}}\{\omega\mathbb{E}_{\bm{x}\in\mathcal{F}[\mathcal{B};\bm{p}]}\bm{D}_{\text{KL}}(\bm{p}||\bm{\tilde{p}})\\ +(1-\omega)\mathbb{E}_{\bm{x}\in\mathcal{F}[\mathcal{B};\bm{\tilde{p}}]}\bm{D}_{\text{KL}}(\bm{\tilde{p}}||\bm{p})\},\end{split}$
where $\omega$ is a random variable drawn from a Bernoulli distribution $\mathcal{B}(0.5)$ (Sun et al., 2022a). Note that the loss gradient is back-propagated through both the perturbed and the original predicted probabilities.
@@
3.4. Adaptive Adjustment Mechanism
Domain adaptation using synthetic data remains a challenging task, especially for long-tailed distributed maritime objects across diverse weather conditions and multiple categories, where substantial variations in sample difficulty exist. Accordingly, we propose an Adaptive Adjustment Mechanism based on curriculum learning and adversarial training, in order to avoid model falling into local optimal, reduce data overfitting and avoid possible model collapse. By imitating the ”step by step” learning process of human, the curriculum learning gradually guides the adversarial training to realize the alignment of source domain and target domain. We make models quickly establish basic representations by firstly feeding easier samples, and then gradually introducing the harder ones. Traditional curriculum learning relied on prior knowledge (usually using fixed values with experience) to evaluate sample difficulty, which leads to lower effectiveness in the later period of training (Wang et al., 2021). Recent method (Jiang et al., 2025) tried using losses as dynamic difficulty evaluation. In this paper, we propose the difficulty score for our task as a combination of prior score and dynamic loss (Eq. LABEL:comprehensive_score) to select source domain samples (i.e., the generated data) for training.
Firstly, we design a prior score as the Image Quality Assessment (IQA) score weighted by weather condition. Specifically, we utilize CLIP IQA (Chen and Mo, 2022) for the prior IQA analysis of the source domain images. According to different weather conditions, the pre-defined weights are:{sunny: 5, sunset & night: 4.5, cloudy: 4, foggy:3.5, rainstorm: 3}. Subsequently, higher image quality and better weather condition contribute to a higher prior score. Secondly, the rest part of difficulty score is based on the standard cross-entropy loss (Eq. 1), where more certain category prediction leads to higher dynamic score. The difficulty score is formulated as follows:
(9) ${difficulty}\_{score}=(1-\phi){prior}\_{score}+\phi(1-\mathcal{L}_{cls}),$
@@ 3.4. Adaptive Adjustment Mechanism
(10) ${A}_{t}=\operatorname{argmax}_{\hat{A}:|\hat{A}|\geq\lambda(t)\cdot|A|}{difficulty}\_{score},$
where $A$ denotes the source domain data, the $|A|$ and $|\hat{A}|$ represent the size of $A$ and $\hat{A}$. $\lambda(t)$ is used to schedule the proportion of source training samples at the t-th epoch, which is defined as:
(11) $\lambda(t)=\lambda_{0}+(1-\lambda_{0})(1-e^{-kt/T}),$
where $\lambda_{0}$ is the initial proportion of source training samples, which is set to 0.5, and $T$ is the total number of epochs. Note that $k$ is set to 2 for the amount of data increasing rapidly in the early stage of training and slowly in the late stage of training, in order to avoid making it difficult to generalize the model to difficult samples due to long-term training of high-score samples.
Furthermore, in the classifier refinement strategy, the selection of appropriate values for the perturbation scalar $\gamma$ and the random offset refinement loss weight $\beta$ is crucial. Excessively large perturbations result in a collapse of the predicted class distribution, whereas small perturbations may fail to fully leverage their potential benefits. Given that the target domain is entirely unlabeled and domain adaptation tasks vary significantly even within the same dataset, it is desirable to adjust these parameters adaptively. At the initial stage of training, the model is not stable, and too much perturbation may lead to convergence difficulties. With the increase of the number of iterations, the proportion of disturbance should be appropriately increased to reduce overfitting. Inspired by SSRT (Sun et al., 2022a), we plan to introduce random perturbations dynamically, by setting an an adaptive scalar.
@@
(12) $\mu(n)=\begin{cases}\sin\left(\frac{\pi}{2N}{n}\right)&\text{if }n<N\\ 1.0&\text{otherwise}\end{cases},$
@@
4. Experiments
4.1. Dataset Construction
@@
AIMO consists of 66,626 images with category labels and multiple weather and illumination conditions. Taking advantage of the strong generation capabilities of Stable Diffusion-WebUI [^1] https://github.com/AUTOMATIC1111/stable-diffusion-webui.git, we create most of the maritime objects using simple text prompts. In order to make the generated data closer to the real data, we utilize techniques such as LoRA, image embeddings (Gal et al., 2022) and hypernetworks (Ha et al., 2016) to assist image generation. For instance, the generated results of warship series, e.g., aircraft carriers, are commonly not satisfactory, so we utilize a LoRA model  [^2] https://www.seaart.ai/models/detail/f8078759a92ae10610a92f8683f893da for better generation quality. The use ratio of LoRA is manually adjusted, according to the generation quality of different categories. In addition, for some categories that are beyond the coverage of the generated model, e.g., barges, tugs and so on, we select a small amount of real maritime objects images for specific categories to make image embeddings by extracting features. The produced image embeddings serve as a prior pseudo-prompt in the generation process, in order to enhance the authenticity and diversity of the generated samples. Considering the complexity of the sea surface environment, we train a hypernetwork on WSODD (Zhou et al., 2021) to achieve sea surface environment style patches, which make the generated data cover a wider range of weather and illumination conditions. As a small neural network, hypernetworks dynamically adjust the weight parameters of the main model (i.e., Stable Diffusion).
RMO consists of 32,418 images collected from multiple maritime object datasets and websites. Specifically, we exclude images that lack the object, as well as those depicting the object in sketch form. And for images containing multiple objects, we crop out each object separately.
@@
4.2. Training Details
We initialize ViT-B/16 (Dosovitskiy et al., 2021) with ImageNet (Russakovsky et al., 2015). We use the SGD optimizer (Le et al., 2011) with a constant learning rate of 0.002 for all our experiments and we set batch size to 16. Our backbone network contains 12 layers of ViT blocks, and we choose blocks 0, 4, and 8 as alternate layers. For features of each batch, a perturbation layer is randomly selected from the alternate layers, and the offsets are respectively added for the source domain and the target domain features in this layer. We set the hyper parameters $\alpha=0.3$, $\beta=1$, $\gamma=0.2$, and $\kappa=0.4$, according to the discussion in Sec 4.6.
Figure 3. Validation experiments with our proposed method and SCAN. (a): Comparison of accuracy on RMO with multiple weather & illumination conditions. (b): Comparison of accuracy on RMO with different categories.
4.3. Effectiveness of AIMO
@@
Firstly, we selected three no-reference image quality assessment metrics (Chen and Mo, 2022), ENTROPY, NIQE, and CLIPIQA to evaluate the quality of AIMO . ENTROPY is used to measure the complexity and information of images. NIQE is used to assess the natural quality of an image. CLIPIQA is used to evaluate image quality by using CLIP. Higher values of these metrics indicate better image quality. Also, we calculated the FID (Chen and Mo, 2022) metric between AIMO and RMO, which characterizes the difference in feature distribution. The results are shown in Table 1. ENTROPY over 7 and CLIPIQA close to 0.8 indicate that AIMO images have rich details, which contain valuable information to transfer. However, the lower NIQE and higher FID indicate a clear difference between AIMO and natural maritime object images, which requires effective domain adaptation.
In Fig.3, we compare our proposed method with SCAN (Van Gansbeke et al., 2020), an unsupervised clustering method that works only on unlabeled RMO. Real-world maritime object images are rarely taken under bad weather or uneven illumination conditions, and have obvious long-tail issues. Therefore, only using real-world data for training easily leads to unbalanced classification results.
@@
Table 1. Image quality assessment on AIMO
@@
Table 2. Comparison with other UDA method
@@
ViT-small (Dosovitskiy et al., 2021) 56.456
ViT-base (Dosovitskiy et al., 2021) 60.693
Swin-B (Liu et al., 2021) 58.761
CDTrans (Xu et al., 2021) 59.183
PMTrans (Zhu et al., 2023) 65.463
TVT (Yang et al., 2023) 62.260
PDA (Bai et al., 2024) 65.891
Ours 73.315
@@
4.4. Effectiveness of Domain Adaptation
We compare our proposed method with other UDA methods, e.g., ViT-small (Dosovitskiy et al., 2021), ViT-base (Dosovitskiy et al., 2021), Swin-B (Liu et al., 2021), CDTrans (Xu et al., 2021), PMTrans (Zhu et al., 2023), TVT (Yang et al., 2023), and PDA (Bai et al., 2024) in Table 2.

### v3

@@ 3.2 Generalization Enhancement Based on CLIP
VLMs exhibit exceptional performance across various distributions, attributed to the vast diversity of distributions seen during their training radford2021learning, which leverage contrastive pre-training methods zhai2022lit for learning the correlation between the embeddings of matching image-text pairs. For instance, CLIP radford2021learning are trained jointly on 400 million image-text pairs, yielding remarkable generalization across different data distributions. Despite the considerable zero-shot generalization of VLMs in classification tasks, effectively adapting them to downstream tasks remains a great challenge. It is expected to leverage the knowledge of VLM in a portable way. Knowledge distillation not only avoids large-scale parameter tuning of VLMs, but also helps to realize lightweight reasoning. For UDA task, it is an effective strategy that enhances cross-domain representation across the open semantic and visual space of VLMs, in order to bridge the domain gap. Compared to a single label, VLMs have more descriptive information, which is beneficial of making full use of the AIMO’s rich label information as semantic consistency constraint. Therefore, we propose a Generalization Enhancement module to efficiently distill the abundant knowledge from CLIP into the adversarial UDA network. We distill the knowledge of text embeddings and image embeddings extracted by frozen VLMs into feature extractor of ViT to enhance feature representations, by carrying out a two-branch constraint to simultaneously align backbone features with semantic information and visual information from VLMs.
@@
Specifically, we perform a Self-Knowledge Distillation on the source domain data based on CLIP. The feature invariance is constructed in the source domain by Self-Knowledge Distillation, which makes the model more robust to the distribution migration in the target domain. Firstly, We utilize the frozen text encoder and image encoder of VLMs to extract embedded features, which update the source domain image representations as a "teacher signal". And then the similarity between text embeddings, image embeddings and backbone features is calculated respectively, as constraints to supervise the training process, in order to realize a double alignment of semantic space and visual space. Given labels for category and weather conditions, as well as images from the source domain as inputs to CLIP, the source domain features of Transformer Blocks are aligned with the text-encoder and image-encoder features of CLIP in the Generalization Enhancement module. As proper text prompt allow for better control of invariance in text embeddings addepalli2024leveraging, we use "A photo of a {class} in {domain}" to represent the semantic embedding. "{domain}" is referred to the weather and illumination condition of the source image. Also, we compared the classification results under various prompt templates in Sec 4.5.
We denote the CLIP’s text embedding for "A photo of a {class} in {domain}" and image embedding of the source domain input $(\bm{x}^{s}_{i},\bm{y}^{s}_{i})$ as $\bm{T}_{{y}^{s}_{i}}$ and $\bm{I}_{\bm{x}^{s}_{i}}$. The generalization enhanced feature corresponding to the input $\bm{x}^{s}_{i}$ are denoted as $\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}_{i}}$. The following Self-Knowledge Distillation loss is used to refine the aligned features:
$\mathcal{L}_{SKD}=-\mathbb{E}_{(\bm{x},\bm{y})\in\mathcal{D}_{s}}\sum_{i=1}^{n_{s}}\{\cos(\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}_{i}},\bm{T}_{{y}^{s}_{i}})+\cos(\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}_{i}},\bm{I}_{\bm{x}^{s}_{i}})\}.$ (5)
@@
3.3 Classifier Refinement
We regularize the latent feature spaces of the transformer backbone through classifier refinement using perturbed source and target domain features to reduce overfitting . Incorporating a random offset into the data is a prevalent data augmentation technique that is widely employed to alleviate overfitting. It may be better to perform data augmentation operations at hidden layers rather than directly manipulating the input images verma2019manifold. The optimal layer for introducing perturbations differs from various tasks. Perturbing relatively deep layers tends to yield superior performance but increases the risk of model collapse, and randomly choosing one transformer block proves to be more robust compared with perturbing any individual layer from them sun2022safe. Therefore, we select a transformer block at random, and then add a random offset to the input token sequence of the chosen block.
@@ 3.3 Classifier Refinement
$\bm{D}_{\text{KL}}(\bm{p}||\bm{\tilde{p}})=\sum_{i=1}\bm{p}(i)\log\frac{\bm{p}(i)}{\bm{\tilde{p}}(i)}.$ (7)
$\bm{D}_{\text{KL}}(\bm{\tilde{p}}||\bm{p})$ is the calculated similarly. And it is more robust to combine the two KL divergence together in random offset loss sun2022safe. To achieve reliable predictions, we define a Confidence Filter $\mathcal{F}$ as
$\mathcal{F}[\mathcal{D};\bm{p}]=\{\bm{x}\in\mathcal{D}|\max(\bm{p})>\kappa\},$ (8)
@@
$\begin{split}\mathcal{L}_{offset}=\mathbb{E}_{\mathcal{B}\subseteq\mathcal{D}_{\bm{s}}\cup\mathcal{D}_{\bm{t}}}\{\omega\mathbb{E}_{\bm{x}\in\mathcal{F}[\mathcal{B};\bm{p}]}\bm{D}_{\text{KL}}(\bm{p}||\bm{\tilde{p}})+(1-\omega)\mathbb{E}_{\bm{x}\in\mathcal{F}[\mathcal{B};\bm{\tilde{p}}]}\bm{D}_{\text{KL}}(\bm{\tilde{p}}||\bm{p})\},\end{split}$ (9)
where $\omega$ is a random variable drawn from a Bernoulli distribution $\mathcal{B}(0.5)$ sun2022safe. Note that the loss gradient is back-propagated through both the perturbed and the original predicted probabilities.
@@
3.4 Adaptive Adjustment Mechanism
Domain adaptation using synthetic data remains a challenging task, especially for long-tailed distributed maritime objects across diverse weather conditions and multiple categories, where substantial variations in sample difficulty exist. Accordingly, we propose an Adaptive Adjustment Mechanism based on curriculum learning and adversarial training, in order to avoid model falling into local optimal, reduce data overfitting and avoid possible model collapse. By imitating the "step by step" learning process of human, the curriculum learning gradually guides the adversarial training to realize the alignment of source domain and target domain. We make models quickly establish basic representations by firstly feeding easier samples, and then gradually introducing the harder ones. Traditional curriculum learning relied on prior knowledge (usually using fixed values with experience) to evaluate sample difficulty, which leads to lower effectiveness in the later period of training wang2021survey. Recent method DBLP:journals/corr/abs-2502-04628 tried using losses as dynamic difficulty evaluation. In this paper, we propose the difficulty score for our task as a combination of prior score and dynamic loss (Eq. 10) to select source domain samples (i.e., the generated data) for training.
Firstly, we design a prior score as the Image Quality Assessment (IQA) score weighted by weather condition. Specifically, we utilize CLIP IQA pyiqa for the prior IQA analysis of the source domain images. According to different weather conditions, the pre-defined weights are:{sunny: 5, sunset & night: 4.5, cloudy: 4, foggy:3.5, rainstorm: 3}. Subsequently, higher image quality and better weather condition contribute to a higher prior score. Secondly, the rest part of difficulty score is based on the standard cross-entropy loss (Eq. 1, where more certain category prediction leads to higher dynamic score. The difficulty score is formulated as follows:
${difficulty}\_{score}=(1-\phi){prior}\_{score}+\phi(1-\mathcal{L}_{ce}),$ (10)
@@ 3.4 Adaptive Adjustment Mechanism
${A}_{t}=\operatorname{argmax}_{\hat{A}:|\hat{A}|\geq\lambda(t)\cdot|A|}{difficulty}\_{score},$ (11)
[…]
$\lambda(t)=\lambda_{0}+(1-\lambda_{0})(1-e^{-kt/T}),$ (12)
where $\lambda_{0}$ is the initial proportion of source training samples, which is set to 0.5, and $T$ is the total number of epochs. Note that $k$ is set to 2 for the amount of data increasing rapidly in the early stage of training and slowly in the late stage of training, in order to avoid making it difficult to generalize the model to difficult samples due to long-term training of high-score samples. Also, the $\tau=5t/T$ in the focal loss varies linearly with $t$ in the range [0, 5], which means that the model loss will gradually pay more attention to difficult samples as the training goes on.
Furthermore, in the classifier refinement strategy, the selection of appropriate values for the perturbation scalar $\gamma$ and the random offset refinement loss weight $\beta$ is crucial. Excessively large perturbations result in a collapse of the predicted class distribution, whereas small perturbations may fail to fully leverage their potential benefits. Given that the target domain is entirely unlabeled and domain adaptation tasks vary significantly even within the same dataset, it is desirable to adjust these parameters adaptively. At the initial stage of training, the model is not stable, and too much perturbation may lead to convergence difficulties. With the increase of the number of iterations, the proportion of disturbance should be appropriately increased to reduce overfitting. Inspired by SSRT sun2022safe, we plan to introduce random perturbations dynamically, by setting an an adaptive scalar.
@@
$\mu(n)=\begin{cases}\sin\left(\frac{\pi}{2N}{n}\right)&\text{if }n<N\\ 1.0&\text{otherwise}\end{cases},$ (13)
@@
4 Experiments
4.1 Dataset Construction
@@
AIMO consists of 66,626 images with category labels and multiple weather and illumination conditions. Taking advantage of the strong generation capabilities of Stable Diffusion-WebUI [^1] https://github.com/AUTOMATIC1111/stable-diffusion-webui.git, we create most of the maritime objects using simple text prompts. In order to make the generated data closer to the real data, we utilize techniques such as LoRA, image embeddings gal2022image and hypernetworks ha2016hypernetworks to assist image generation. For instance, the generated results of warship series, e.g., aircraft carriers, are commonly not satisfactory, so we utilize a LoRA model  [^2] https://www.seaart.ai/models/detail/f8078759a92ae10610a92f8683f893da for better generation quality. The use ratio of LoRA is manually adjusted, according to the generation quality of different categories. In addition, for some categories that are beyond the coverage of the generated model, e.g., barges, tugs and so on, we select a small amount of real maritime objects images for specific categories to make image embeddings by extracting features. The produced image embeddings serve as a prior pseudo-prompt in the generation process, in order to enhance the authenticity and diversity of the generated samples. Considering the complexity of the sea surface environment, we train a hypernetwork on WSODD zhou2021image to achieve sea surface environment style patches, which make the generated data cover a wider range of weather and illumination conditions. As a small neural network, hypernetworks dynamically adjust the weight parameters of the main model (i.e., Stable Diffusion).
RMO consists of 32,418 images collected from multiple maritime object datasets and websites. The main sources of RMO are: Marvel gundogdu2017marvel, Mcships zheng2020mcships, ABOships iancu2021aboships, MAR-ships bloisi2015argos, CIB-ships meng2022fine, WSODD zhou2021image, DVTR zhang2024viewpoint, Ships Image Dataset [^3] https://www.kaggle.com/datasets/vinayakshanawad/ships-dataset, Boat types recognition [^4] https://www.kaggle.com/datasets/clorichel/boat-types-recognition, Buoyandboat Computer Vision Project [^5] https://universe.roboflow.com/univeristy-of-southeastern-norway/buoyandboat. Then we filter these data to form the RMO. Specifically, we exclude images that lack the object, as well as those depicting the object in sketch form. And for images containing multiple objects, we crop out each object separately.
@@
4.2 Training Details
We initialize ViT-B/16 DBLP:conf/iclr/DosovitskiyB0WZ21 with ImageNet russakovsky2015imagenet. We use the SGD optimizer le2011optimization with a constant learning rate of 0.002 for all our experiments and we set batch size to 16. Our backbone network contains 12 layers of ViT blocks, and we choose blocks 0, 4, and 8 as alternate layers. For features of each batch, a perturbation layer is randomly selected from the alternate layers, and the offsets are respectively added for the source domain and the target domain features in this layer. We set the hyper parameters $\alpha=0.3$, $\beta=1$, $\gamma=0.2$, and $\kappa=0.4$, according to the discussion in Sec 4.6.
4.3 Effectiveness of AIMO
@@
Firstly, we selected three no-reference image quality assessment metrics pyiqa, ENTROPY, NIQE, and CLIPIQA to evaluate the quality of AIMO . ENTROPY is used to measure the complexity and information of images. NIQE is used to assess the natural quality of an image. CLIPIQA is used to evaluate image quality by using CLIP. Higher values of these metrics indicate better image quality. Also, we calculated the FID pyiqa metric between AIMO and RMO, which characterizes the difference in feature distribution. The results are shown in Table 1. ENTROPY over 7 and CLIPIQA close to 0.8 indicate that AIMO images have rich details, which contain valuable information to transfer. However, the lower NIQE and higher FID indicate a clear difference between AIMO and natural maritime object images, which requires effective domain adaptation.
In Fig.3, we compare our proposed method with SCAN van2020scan, an unsupervised clustering method that works only on unlabeled RMO. Real-world maritime object images are rarely taken under bad weather or uneven illumination conditions, and have obvious long-tail issues. Therefore, only using real-world data for training easily leads to unbalanced classification results.
@@
Table 1: Image quality assessment on AIMO
@@
Table 2: Comparison with other UDA methods on RMO
@@
ViT-small DBLP:conf/iclr/DosovitskiyB0WZ21 56.456
ViT-base DBLP:conf/iclr/DosovitskiyB0WZ21 60.693
Swin-B liu2021swin 58.761
CDTrans xu2021cdtrans 59.183
PMTrans zhu2023patch 65.463
TVT yang2023tvt 62.260
PDA bai2024prompt 65.891
UniMoS li2024split 69.798
DAMP du2024domain 72.398
Ours 73.675
Table 3: Comparison with other UDA methods on Mini-DomainNet
Method Acc %
ViT-base DBLP:conf/iclr/DosovitskiyB0WZ21 65.7
CLIP radford2021learning 82.8
DAPromp DBLP:journals/tnn/GeHXLSLH25 85.8
ADCLIP DBLP:conf/iccvw/SinghaPJB23 86.9
UniMoS li2024split 87.3
DAMP du2024domain 87.6
Ours 87.5
@@
4.4 Effectiveness of Domain Adaptation
We compare our proposed method with other UDA methods, e.g., ViT-small DBLP:conf/iclr/DosovitskiyB0WZ21, ViT-base DBLP:conf/iclr/DosovitskiyB0WZ21, Swin-B liu2021swin, CDTrans xu2021cdtrans, PMTrans zhu2023patch, TVT yang2023tvt, PDA bai2024prompt, UniMoS li2024split, and DAMP du2024domain in Table 2. In summary, our proposed method performs better than other recent methods, and obtains an average classification accuracy of 73.675% on RMO.